# Entanglement scaling behaviors of free fermions on hyperbolic lattices

### v1

@@ IV Entanglement entropy scaling of gapped free-fermion systems
where $c^{\dagger}_{s(p),i(j)}$ represents fermionic creation operator at the $s(p)$-orbital of site $i(j)$. $t_{1}$ and $t_{2}$ are hopping amplitudes. We can still use Haydock recursion method to compute DOS and verify that $H_{2}$ is gapped as we did in Sec III. For instance,in Fig. 7(a), we show the DOS of Hamiltonian $H_{2}$ with $t_{1}=1$ and $t_{2}=1$ on $\{4,5\}$ lattice, which lead to a gapped region $[-1,1]$.
@@ V Discussions
In this paper, we have numerically studied the scaling behavior of entanglement entropy of gapped free fermions as well as gapless free fermions with finite DOS on hyperbolic lattice. We find that for both gapped and gapless systems, the EE scales according to a rigorous area law scaling $S_{A}=aL_{A}+\cdots$. Although the gapped case fulfills our expectation in Euclidean geometry, the super-area law in gapless systems breaks down in contrast. Additionally, the scaling coefficient of area law in gapless systems is positively related to the DOS. This scaling behavior of EE is unique in hyperbolic geometry. On Euclidean lattice, the super-area law of gapless free fermions with finite DOS demonstrates that the entanglement is enhanced by the fermionic statistics and the quantum correlation of the infinite fermion modes near the Fermi surface [12, 14]. Compared to Euclidean case, the area law of gapless systems on hyperbolic lattices suggests the exotic properties induced by the hyperbolic geometry and the generalization of Widom conjecture to hyperbolic lattices. The generalized Bloch theory, or hyperbolic band theory (HBT), of hyperbolic lattice could provide a description through representation theory for non-Abelian translation group of hyperbolic lattice, thereby describing the generalized momentun space and may yield a momentun space formula for EE [43, 42, 44].
Moreover, we find that the EE of a free-fermion system behaves differently under different geometries, reflecting the nature of the spatial geometry where the system is embedded, e.g., fractal lattice discussed in Ref. [45]. As it is feasible to simulate entanglement experimentally [52] while the photonic and circuit experimental realization of hyperbolic lattice has been achieved [27, 26, 30, 28, 29, 31], this may provide us with a novel approach to study the geometry of the quantum system through entanglement. Additionally, the area law of EE in both gapped and gapless systems suggests that it is efficient to study correlated systems on hyperbolic lattices with gapless emergent fermions by tensor-network-type numerical techniques [53, 18, 54]. We hope that our work can provide some inspiration to related fields in the future. Another interesting future direction is to study entanglement of non-Hermitian systems [69, 70, 64, 65, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87] on hyperbolic lattice, which is much more practical in, e.g., phononic systems where gain and loss are natural.
@@ Acknowledgements.
This work was supported by National Natural Science Foundation of China (NSFC) Grant No. 12074438 and Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices under Grant No. 2022B1212010008.
@@ Appendix A Hyperbolic lattice
In this section we give details of constructing hyperbolic lattices discussed in Sec. II.
@@ C.2 Numerical results of DOS
Appendix D Asymptotic behavior of scaling coefficient of area law

### v2

@@ IV Entanglement entropy scaling of gapped free-fermion systems
where $c^{\dagger}_{s(p),i(j)}$ represents fermionic creation operator at the $s(p)$-orbital of site $i(j)$. $t_{1}$ and $t_{2}$ are hopping amplitudes. We can still use Haydock recursion method to compute DOS and verify that $H_{2}$ is gapped as we did in Sec. III. For instance,in Fig. 7(a), we show the DOS of Hamiltonian $H_{2}$ with $t_{1}=1$ and $t_{2}=1$ on $\{4,5\}$ lattice, which lead to a gapped region $[-1,1]$.
@@ V Discussions
In this paper, we have numerically studied the scaling behavior of entanglement entropy of gapped free fermions as well as gapless free fermions with finite DOS on hyperbolic lattice. We find that for both gapped and gapless systems, the EE scales according to a rigorous area law scaling $S_{A}=aL_{A}+\cdots$. Although the gapped case fulfills our expectation in Euclidean geometry, the super-area law in gapless systems breaks down in contrast. Additionally, the scaling coefficient of area law in gapless systems is positively correlated to the DOS. This scaling behavior of EE is unique in hyperbolic geometry. On Euclidean lattice, the super-area law of gapless free fermions with finite DOS demonstrates that the entanglement is enhanced by the fermionic statistics and the quantum correlation of the infinite fermion modes near the Fermi surface [12, 14]. These observations show that the underlying geometry can significantly influence the entanglement behavior of ground states of free-fermion systems, similar to our previous findings in fractal geometry [15], as summarized in Table. 1. Compared to Euclidean case, the area law reveals the presence of exotic properties of fermions on hyperbolic lattice. The further study of this area law might raise the question for a generalized conjecture of correlation matrices with symmetry of hyperbolic lattice, while the possibility of generalization of Swingle’s argument for EE to the hyperbolic case through HBT also merits future research [56, 55, 57].
Notably, many studies suggest the relationship between the entanglement and the geometry of AdS space, particularly in the context of hyperbolic lattice [34, 36, 37, 38]. Ref. [34] experimentally studies AdS/CFT correspondence, confirming the EE for entanglement wedge as subsystem of bulk weakly-coupled scalar field is consistent with the RT formula [19]. Inspired by the experimental progress, the experimental simulation of Gaussian fermionic field theory to study EE is worthy of further studying. Additionally, numerical studies of spin models also point out the non-trivial behaviors of correlation functions and entanglement of spin models on hyperbolic lattice [36, 37, 38]. Based on our numerical results and these works, along with the theoretical research on the relationship between EE and AdS space geometry, the investigation within the framework of field theory and holography to understand the area-law EE of ground states for free-fermion systems on hyperbolic lattice is an interesting topic for future work [17, 18, 19, 20, 21, 22, 23].
As it is feasible to simulate entanglement experimentally [66] while the experimental simulation of hyperbolic lattice has been achieved [30, 31, 32, 33, 34, 35], this may provide us with a novel approach to study the geometry of the quantum system through entanglement. Furthermore, the area law of EE in both gapped and gapless systems suggests that it is efficient to study correlated systems on hyperbolic lattices with gapless emergent fermions by tensor-network-type numerical techniques [67, 25, 68]. We hope that our work can provide some inspiration to related fields in the future. Another interesting future direction is to study entanglement of non-Hermitian systems [82, 83, 77, 78, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100] on hyperbolic lattice, which is much more practical in, e.g., phononic systems where gain and loss are natural.
@@ Acknowledgements.
This work was in part supported by National Natural Science Foundation of China (NSFC) Grant No. 12474149 and No. 12074438. The calculations reported were performed on resources provided by the Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices, No. 2022B1212010008.
@@ Appendix A Hyperbolic lattice
In this section we give details of constructing hyperbolic lattices discussed in Sec. II. Additionally, we discuss the geometrical properties of hyperbolic lattice as well as the volume law of EE.
@@ A.2 Exponential growth of the size of the hyperbolic lattice
A.3 Entanglement volume law on hyperbolic lattice
In this section, we show that $L_{A}$ of a subsystem approaches a finite fraction of the total number of lattice sites in the subsystem $N(V_{A})$ in the thermodynamical limit $n\rightarrow\infty$, and discuss the volume law of EE on hyperbolic lattice.
Consider a $\{p,q,n\}$ lattice as a subsystem $A$ in a larger lattice. The $L_{A}$ defined in Sec. II.2 can be expressed as:
$L_{A}=(q-2)N_{n,2}+(q-3)N_{n,3}\,.$ (18)
By solving for $N_{n,2}$ and $N_{n,3}$ similar to the approach in Appendix A.2, we find that the leading-order terms of $N_{n,2}$ and $N_{n,3}$ coincide with those of $N_{n}$. Therefore, $L_{A}/N$ becomes a finite fraction for sufficiently large subsystems or in the thermodynamical limit $n\rightarrow\infty$, e.g., approximately $1.732$ for the $\{4,5\}$ lattice.
The area law indicates that the EE is proportional to the degrees of freedom $dN(L^{D-1}_{A})$ on the boundary of the subsystem, where $d$ is the local Hilbert space on a lattice site, $D$ is the spacial dimension and $N(L^{D-1}_{A})$ is the number of boundary lattice sites of the subsystem. On Euclidean geometry, $N(L^{D-1}_{A})$ is proportional to the boundary area $L^{D-1}_{A}$, where $L_{A}$ is the linear size of the subsystem [7]. Meanwhile, $N(L^{D-1}_{A})$ is equal to the total number of lattice bonds connecting the lattice sites in two complementary subsystems. Therefore, the scaling $S_{A}\sim L^{D-1}_{A}$ is referred to as the area law. In our work presented here, we also adopt the total number of bonds as the linear size $L_{A}$, as clarified in Sec. III.2, and numerically find the scaling $S_{A}\sim L_{A}$ still holds for both gapless (with finite DOS) and gapped free-fermion systems on hyperbolic lattices. Hence, we refer to this scaling on hyperbolic lattices as the area law.
The volume law indicates that the EE is proportional to the total degrees of freedoms $dN(V_{A})$ within the subsystem, where $N(V_{A})$ is the total number of lattice sites in the subsystem. In Euclidean geometry, $N(V_{A})$ is proportional to the volume $V_{A}\sim L^{D}_{A}$ of the subsystem. Therefore, the scaling $S_{A}\sim L^{D}_{A}$ is referred to as the volume law. On hyperbolic lattices, however, the area of a region can scale as a finite fraction of its volume in the asymptotic limit (i.e., for sufficiently large subsystems), leading to $V_{A}\sim L_{A}$ geometrically. Consequently, on hyperbolic lattices, the area law can also be interpreted as $S_{A}\sim L_{A}\sim V_{A}$, which may alternatively be referred to as the volume law. However, to ensure the definition of area law is consistent on both Euclidean and hyperbolic lattices, we still regard the scaling $S_{A}\sim L_{A}$ on hyperbolic lattices as the area law.
@@ C.2 Numerical results of DOS
Figure 11: Finite-size scaling analysis of EE by the example of $\{8,3\}$ lattice. (a) EE for the same subsystems computed on $\{8,3,n\}$ lattice with $n=4,5,6$. (b) EE for two specific subsystems ($V_{A}=48$ and $200$ sites) computed on $\{8,3,n\}$ lattice, where dashed lines show ±$\pm 5\%$ intervals of EE computed on $\{8,3,6\}$ lattice. Results are computed for Hamiltonian $H_{1}$ with $t=1$ and $\mu=0$.
Table 3: Results of adjusted coefficient of determination $\bar{R}^{2}$ and dimensionless error estimations, including mean absolute percentage error (MAPE) and relative absolute error (RAE) for EE data shown in Fig. (4) are presented. In Euclidean case, the super-area law $S_{A}/\log L_{A}=a_{2}L_{A}+b_{2}$ is better fit, while in hyperbolic case the area law $S_{A}=a_{1}L_{A}+b_{1}$ is better.
[…]
$\{4,4,40\}$ area 0.998466 3.66006 $3.69134\times 10^{-2}$
super-area 0.999929 0.338284 $7.64020\times 10^{-3}$
$\{3,7,9\}$ area 0.999962 0.495449 $5.79884\times 10^{-3}$
super-area 0.998026 3.12212 $4.06104\times 10^{-2}$
$\{4,5,6\}$ area 0.999918 0.832573 $8.48647\times 10^{-3}$
super-area 0.998041 2.55588 $4.20934\times 10^{-2}$
Figure 12: Column (a-c) correspond to the EE data of $\{4,4,40\}$, $\{3,7,9\}$, $\{4,5,6\}$ shown in Fig. (4). In row (1-2), we present the linear fits $S_{A}=a_{1}L_{A}+b_{1}$ and $S_{A}/\log L_{A}=a_{2}L_{A}+b_{2}$, along with the adjusted coefficient of determination $\bar{R}^{2}$. The blue lines represent the respective linear fit functions. In row (3), we plot $S_{A}/L_{A}\log L_{A}$ as a function of $L_{A}$, where dashed lines represent ±$\pm 5\%$ intervals of the data point corresponding to the biggest $L_{A}$.
Appendix D Finite-size scaling analysis of EE
As we focus on EE for subsystems in the bulk, in this section we perform the finite-size scaling analysis to demonstrate that the boundary effect is considered in our numerical computations. We take the $\{8,3\}$ lattice for Hamiltonian $H_{1}$ with $t=1$ and $\mu=0$ as an example. In Fig. 11(a), we show the EE for the same subsystems (partition i is taken) computed on $\{8,3,4\}$ ($768$ sites), $\{8,3,5\}$ ($2888$ sites) and $\{8,3,6\}$ ($10800$ sites) lattice respectively, where the largest subsystem studied here is identical to $\{8,3,3\}$ lattice ($V_{A}=200$ sites). The optimal fit is obtained when $\alpha$ in the fitting function $S_{A}=aL_{A}^{\alpha}+b$ equals to $0.859$, $1.010$ and $0.997$ for $\{8,3,4\}$, $\{8,3,5\}$ and $\{8,3,6\}$ lattice respectively. In Fig. 11(b), two specific subsystems, identical to $\{8,3,2\}$ ($V_{A}=48$ sites) and $\{8,3,3\}$ ($V_{A}=200$ sites) lattice respectively, are chosen to compute EE on $\{8,3,n\}$ lattices with $n=3,4,5,6$. Notably, the $\{8,3,3\}$ subsystem is not computed on $\{8,3,3\}$ lattice itself. The gray dashed lines represent ±$\pm 5\%$ intervals of the EE computed on $\{8,3,6\}$ lattice. From these results, we observe that by minimizing boundary effects through enlarging the lattice, we obtain a linear fit of the EE for subsystems in the bulk. Through the finite-size scaling analysis, we show that the boundary effect is considerably excluded to obtain the EE for subsystems in the bulk of the hyperbolic lattice.
Appendix E Numerical analysis of super-area law scaling
In this section, we present the numerical analysis to show that the EE does not scale as super-area law scaling. Specifically, we take the EE data for the gapless free fermions with finite DOS presented in Fig. (4), Sec. III.2 as an example. We anticipate the scaling functions $S_{A}=a_{1}L_{A}+b_{1}$ and $S_{A}/\log L_{A}=a_{2}L_{A}+b_{2}$ and fit the EE data. As logarithms can be hard to detect, we select several dimensionless evaluation metrics to compare these two fits, including the adjusted coefficient of determination $\bar{R}^{2}$, mean absolute percentage error (MAPE) given by $\frac{100\%}{N}\sum_{i}^{N}{\left|\frac{\hat{y}_{i}-y_{i}}{y_{i}}\right|}$, and relative absolute error (RAE) given by $\frac{\sum_{i}^{N}{\left|y_{i}-\hat{y}_{i}\right|}}{\sum_{i}^{N}{\left|y_{i}-\overline{y}\right|}}$. Here $N$ is the total number of data points, $y_{i}$ denotes the true values, $\overline{y}$ denotes the mean of $y$ and $\hat{y}_{i}$ denotes the predicted values. A higher $\bar{R}^{2}$ approaching $1$ and lower error estimations approaching 00 indicate a better fitting. The results of these evaluations are listed in Table. (3), where one can find that area-law fitting is better in hyperbolic case. We also visualize the fittings in row (1-2) of Fig. (12). In row (3), we plot $S_{A}/L_{A}\log L_{A}$ as a function of $L_{A}$ which only converge in the Euclidean case. Therefore, we conclude that the EE does not scale as super-area law.
@@ Appendix E Numerical analysis of super-area law scaling
Appendix F Asymptotic behavior of scaling coefficient of area law